### v3

# Reproducible Machine Learning-based Voice Pathology Detection: Introducing the Pitch Difference Feature

Abstract
Purpose: We introduce a novel methodology for voice pathology detection using the publicly available Saarbrücken Voice Database (SVD) and a robust feature set combining commonly used acoustic handcrafted features with two novel ones: pitch difference (relative variation in fundamental frequency) and NaN feature (failed fundamental frequency estimation).
Methods: We evaluate six machine learning (ML) algorithms — support vector machine, k-nearest neighbors, naive Bayes, decision tree, random forest, and AdaBoost — using grid search for feasible hyperparameters and 20480 different feature subsets. Top 1000 classification models – feature subset combinations for each ML algorithm are validated with repeated stratified cross-validation. To address class imbalance, we apply K-Means SMOTE to augment the training data.
Results: Our approach achieves 85.61%, 84.69% and 85.22% unweighted average recall (UAR) for females, males and combined results respectively. We intentionally omit accuracy as it is a highly biased metric for imbalanced data.
Conclusion: Our study demonstrates that by following the proposed methodology and feature engineering, there is a potential in detection of various voice pathologies using ML models applied to the simplest vocal task, a sustained utterance of the vowel /a:/. To enable easier use of our methodology and to support our claims, we provide a publicly available GitHub repository with DOI 10.5281/zenodo.13771573. Finally, we provide a REFORMS checklist to enhance readability, reproducibility and justification of our approach.
keywords:
Voice pathology detection , voice disorder detection , Saarbrücken Voice Database , SVD , machine learning , REFORMS
journal: Journal of Voice
[…]
[1]organization=Department of Mathematics, Informatics, and Cybernetics, University of Chemistry and Technology, Prague, addressline=Technická 5, city=Prague, postcode=166 28, country=Czech Republic
[…]
[2]organization=Department of Radiological Imaging and Informatics, Tohoku University Graduate School of Medicine, addressline=2-1-1 Katahira, Aoba-ku, city=Sendai, postcode=980-8577, country=Japan
[…]
[3]organization=Department of Mathematics and Physics, University of Campania ”Luigi Vanvitelli”, addressline=Viale Abramo Lincoln 5, city=Caserta, postcode=81100, country=Italy
[…]
[4]organization=Department of Otorhinolaryngology, Faculty Hospital Královské Vinohrady, addressline=Šrobárova 1150/50, city=Prague, postcode=100 34, country=Czech Republic
{graphicalabstract}[Uncaptioned image]
{highlights}
Introducing reproducible machine-learning-based system for voice pathology detection.
Novel pitch difference and NaN features improving classification performance.
Addressing class imbalance with k-means SMOTE, boosting minority class predictions.
Avoiding data leakage by managing multiple recordings per patient effectively.
Performing extensive grid search for optimal features and hyperparameters.
1 Introduction
Voice and speech are fundamental aspects of human communication, playing a crucial role in social interaction, emotional expression, and professional performance. Disorders affecting these functions can have a profound impact on an individual’s quality of life, leading to reduced intelligibility, impaired social interactions, and psychological distress. Voice and speech pathologies encompass a diverse range of conditions, each characterized by distinct alterations in the acoustic and articulatory properties of speech signals. A precise understanding of these pathologies is essential for accurate diagnosis, effective therapeutic intervention, and improved clinical outcomes. [1]
Traditionally, voice disorders have often been described primarily in terms of voice quality disturbances, as seen in conditions such as dysphonia, which is commonly associated with hoarseness, breathiness, or strain. However, this perspective is too narrow to encompass the full spectrum of speech and voice disorders. For instance, dysarthria, a motor speech disorder resulting from neurological impairment, affects articulation, prosody, resonance, and speech duration, rather than merely altering voice quality. Similarly, conditions such as apraxia of speech and neurological voice disorders, including spasmodic dysphonia and hypokinetic dysarthria due to the Parkinson’s disease, lead to a combination of deficits in phonation, speech timing, and intonation. [2, 3] Moreover, emerging research has emphasized the importance of objective acoustic and articulatory analyses in the evaluation of voice and speech disorders.
Advancements in computerized speech analysis, laryngeal imaging, and neurophysiological assessments have provided deeper insights into the pathophysiological mechanisms underlying these conditions. These tools enable clinicians and researchers to characterize speech disorders more precisely, facilitating early diagnosis and the development of targeted interventions.
This involves the processing and analysis of various acoustic features of the voice signal, which can reveal changes or alterations in voice quality caused by specific diseases [4]. In this context, artificial intelligence can be a valid and powerful tool as it may support the decision making process in clinical settings. However, utilization of these solutions imposes many additional tasks, ranging from data collection and preprocessing to ground truth labeling and correct algorithm selection.
Moreover, the choice of appropriate acoustic features is fundamental. The use of acoustic features to characterize pathological voice quality has been investigated in a variety of contexts and for a variety of purposes [5, 6]. These features can provide a quantitative method of assessing voice characteristics that are otherwise difficult to measure. However, there is no standardized set of acoustic measures, making the selection of appropriate acoustic metrics and their interpretation an ongoing challenge.
In this paper on voice pathology detection, we introduce a methodology for Saarbrücken Voice Database (SVD) that effectively addresses the challenge of multiple recordings from the same patients. Without proper handling, such data redundancy can lead to data leakage, an undesirable effect in which the classification model obtains information about the validation set during the training phase, which might lead to overly optimistic classification results [7]. We consider this an important contribution, as this issue, in the context of voice pathology detection, is rarely discussed in the literature.
Additionally, we present two novel features — pitch difference and NaN feature. Several used features are based on the fundamental frequency (f0). The NaN feature reflects the fact, that we were not able to extract f0 values in the analyzed speech signal. The pitch difference feature quantifies the variability of f0 by measuring fluctuations within a single sustained vowel production. We assume that this variability is often altered in pathological voices due to impairments in vocal fold function, making it a possibly useful feature for voice pathology detection. We experimentally verify proposed features’ usefulness in voice pathology detection.
Based on our experimental findings, we emphasize the necessity of training models separately for male and female patients. Given that we do not investigate the causality or direct correlation of individual features with specific pathologies — and instead work with a large feature set — we leverage machine learning (ML) algorithms. Our feature selection process reveals that the optimal feature sets may vary depending on both the sex and ML algorithm. To identify robust feature sets, we train multiple classification models across different combinations of feature subsets (see Table 2).
By computing the mean Matthews correlation coefficient (MCC), we identify the most promising feature – classification model combinations. We then perform repeated stratified 10-fold cross-validation for the top-performing models to estimate the average performance metrics along with their standard deviations, providing insight into the confidence of our results.
To our best knowledge, our work is the first in the field of voice pathology detection to provide fully reproducible results. The code used for computations and feature extraction is available in our repository, ensuring complete transparency and allowing anyone to verify, build upon, or extend our work. By prioritizing reproducibility, we aim to set a standard for rigorous and open research.
2 Related Works
In this chapter, we outline the original contribution related to voice pathology detection present in the literature, limiting our research to works connected to SVD. This choice is due to its comprehensiveness: it is the only voice recording database representing the common voice pathologies. See Section 3.1 for detailed information.
Harar et al. [8] test various acoustic features, such as pitch, jitter, shimmer, harmonic-to-noise ratio (HNR), detrended fluctuation analysis parameters, glottis quotients (open, closed), glottal-to-noise excitation ratio, Teager–Kaiser energy operator, modulation energy, and normalized noise energy as well as mel-frequency cepstral coefficients (MFCC). Moreover, they considered the sound samples and their spectrograms as input, all in combination with XGBoost, IsolationForest, and DenseNet models to determine the pathologic samples, reaching an F1 score of 73.3% and unweighted average recall (UAR) we computed as 73.3%. While, in Gupta et al.[9], features derived from self-supervised learning models, Data2Vec and Wav2Vec, alongside MFCC are explored. The reliability of these features to evaluate voice quality is tested using support vector machine (SVM) and deep neural networks (DNN), achieving an accuracy of 77.83% and UAR we computed as 77.86%.
A different approach is taken by Verde et al.[10], where the authors transform sound wave data into spectrograms and treat classification as an image recognition problem using a convolutional neural network (CNN), achieving 73.93% accuracy and UAR we computed as 70.68%.
In Kotarba & Kotarba[11], the authors use MFCC and gammatone frequency cepstral coefficients (GFCC) with a classification model based on a neural network (NN), achieving 81.84% accuracy. Unlike other studies that use the /a:/ sound for feature extraction, they employ whole sentences. Unfortunately, we were not able to compute UAR, because the reported results appear inconsistent, making it impossible to calculate the metric. Another study by Harar et al.[12] tests a DNN with convolutional layers, using the voice signal as a feature set for the convolutional layer, achieving 68.08% accuracy and UAR we computed as 72.32%. Park et al.[13] combine CNN-based feature extractors with various glsml algorithms, such as SVM and DNN s, achieving an UAR score of 84.97% on the entire SVD dataset.
Finally, Omeroglu et al.[14] test a combination of SVM and CNN-based feature extractors on spectrograms of sound and electroglottogram (EGG) signals. They integrate these with traditional acoustic features like MFCC, linear predictive coefficients (LPC), and f0, reaching an accuracy of 90.10% and UAR we computed as 88.75%. Similarly, they reach an accuracy of 87.41% by extracting the mentioned features and using SVM for classification. However, they reach such high performance without specifying how they handled multiple recordings from the same patient which occur in the dataset, potentially introducing data leakage and therefore reporting an overly optimistic performance.
In another study by Verde et al.[15], the authors take age and sex information from the database and extract various acoustic features from the time domain, such as f0, jitter, shimmer, and HNR. They employ ML algorithms, including boosted trees, SVM, decision trees (DT), naive Bayes (NB), and k-nearest neighbors (KNN), achieving the highest accuracy of 84.5% with a boosted tree model on an imbalanced dataset and UAR we computed as 84.55%. In the follow-up study [16], they expand their work by incorporating MFCC and their derivatives, and test additional ML algorithms such as logistic model tree and instance-based learning algorithms, reaching 85.77% accuracy with a SVM model and UAR we computed as 85.77%. However, they utilized a balanced subset of SVD.
Additionally, among the researched works, there were several works that took a subset of the SVD based on selected pathologies.
Tirronen et al.[17] extract features based on MFCC and self-supervised algorithms from samples of healthy individuals and patients with hyperfunctional dysphonia and vocal fold paresis. Using SVM, they achieve 75.65% and 74.50% accuracy for male and female patients, respectively. Compared to that, Yagnavajjula et al.[18] focus on developing classification models to distinguish between healthy subjects and patients suffering from spasmodic dysphonia and laryngeal nerve paralysis. Their exploration of multi-modal classification methods results in an accuracy of 68.11%.
Further expanding the scope, Junior et al.[19] investigate multi-modal classification for patients with various conditions, including dysphonia, laryngitis, Reinke’s edema, vox senilis, and central laryngeal motion disorder, using energy, zero-crossing rate, and entropy as features. By employing SVM and NN, they achieve an average accuracy of 88.46%.
Similarly, Fan et al.[20] explore multimodal classification for conditions such as nodules, polyps, edema, and paralysis, as well as binary classification between healthy and unhealthy individuals. They utilize MFCC as features and a synthetic minority over-sampling technique (SMOTE)-based method for balancing datasets and reach a maximum F1-score of 90% with a CNN-based model for binary classification and we computed UAR as 90%. However, they do not describe how they handle duplicities in SVD dataset and very likely introduce the data leakage due their methodology.
Ding et al.[21] propose their own approach by combining MFCC and log-mel-frequency spectral coefficients with deep models, using recordings from the SVD and their own database. This approach results in an accuracy of up to 81.6%. Meanwhile, Guedes et al.[22] focus on patients with dysphonia, chronic laryngitis, and vocal cord paralysis, employing features extracted by a VGGish model in combination with a long short-term memory (LSTM) network. They achieve an F1-score of up to 80% in distinguishing between healthy individuals and those with paralysis.
Hemmerling[23] tests quantitative voice parameters combined with a multi-layered perceptron (MLP) model to classify healthy individuals and those with hyperfunctional dysphonia, laryngitis, and recurrent laryngeal nerve paralysis, achieving 87.5% accuracy.
Additionally, AL-Dhief et al.[24] select 280 samples for pathology detection, extracting MFCC features and using them in a DT model, which results in an accuracy of 67.9%.
Expanding the feature set, AnilKumar & Reddy[25] use MFCC, first and second derivatives of MFCC, linear prediction cepstral coefficients, and constant - Q cepstral coefficients with a Bi-LSTM to classify eight selected pathologies, namely dysody, dysphonia, functional dysphonia, hyperfunctional dysphonia, hypofunctional dysphonia, spasmodic dysphonia, vocal cyst polyp, and healthy individuals, achieving an accuracy of 92.7%, which is the only metric they provide. Finally, Tirronen et al.[26] examine the impact of MFCC extraction window length on detecting pathologies in patients with dysphonia and reflux laryngitis. Using SVM, they achieve up to 75.1% accuracy and UAR we computed as 75%.
Finally, multiple works study the use of end-to-end models, which are not solely dependent on hand-crafted features. [27, 28] Liu et al. [29] propose an end-to-end deep learning model for classification of laryngitis and hyperfunctional dysphonia using stacked vowels, while reaching UAR of 72%. Reddy et al. [30] apply a wavelet scattering network to extract features from recordings in combination with MLP for classification. With their end-to-end approach, they reach 81.32% UAR using recordings of sentences as their input.
3 Materials & Methods
To make our methodology more comprehensive, we provide a flowchart to show the process of feature extraction, training and results validation in Figure 1, along with the reference to the relevant tables and figures in the manuscript.
Figure 1: Flow diagram of the proposed methodology
3.1 Data
Voice pathology detection studies often use the MEEI [31], VOICED [32, 33], FEMH Voice Data Challenge 2018 [34], AVPD [35] datasets, and several works use their own datasets. However, the MEEI dataset is no longer available, and the FEMH Voice Data Challenge 2018 is not publicly available. We do not consider the AVPD dataset feasible for our study, as it includes only five types of pathologies: vocal fold cysts, nodules, paralysis, polyps, and sulcus. Thus, it does not reflect the number of various pathologies presented in the general population. The VOICED database contains 3 different pathologies, comprising 21 various disorders, but contains a limited sample size of 208 recordings.
Therefore, we use the SVD [36] for our work as it is publicly available, contains a wide range of pathologies, and is the largest of the databases. The SVD was developed by the Phonetics group at the Department of Language Science and Technology, Saarland University, and is available at https://stimmdb.coli.uni-saarland.de.
Figure 2: Healthy female subject ID 1
Figure 3: Female subject ID 1, suffering with functional dysphonia
The dataset contains data from 1853 patients and includes voice and EGG recordings of various vowels in different pitches, as well as from a German phrase, possibly from multiple recording sessions. For each recording session, information concerning the sex and age of the speaker is also provided, as well as a list of pathologies and diagnoses and any comments.
For our research, we use the recordings containing the sound /a:/ in neutral pitch for pathology detection due to its frequent use in previous studies and clinical protocols [37, 38]. Figure 2 and Figure 3 present examples of healthy and pathological spectrograms, illustrating the /a:/ sound in neutral pitch from a healthy female subject (patient ID 1) and the same subject with functional dysphonia. We fully recognize that many voice as well as speech pathologies go beyond phonatory function and involve articulatory, prosodic, and temporal characteristics, which are more effectively captured in continuous speech.
For feature extraction, we consider some limitations to ensure robust and unbiased results in voice pathology detection. The database contains samples from underage patients. Due to developmental changes at a young age, we exclude any recordings of patients under 18 years of age. Research [39, 40] shows that during development, the characteristics of young voices, such as the fundamental frequency, are distinct from those of fully developed voices. These developmental changes could cause problems for the classification model and reduce its ability to accurately detect pathology. By excluding these age groups, we aim to maintain a more consistent and reliable dataset for our analysis.
Another significant issue is the presence of multiple recordings for some patients. For example, the patient with the talker ID 2027 has 24 recordings of the sound /a:/ in neutral pitch. This repetition poses a risk of data leakage, which can affect the results if not addressed properly.
To prevent this risk in a deterministic way, we select the oldest sample of each type by date and recording ID, resulting in a maximum of two recordings per patient, one healthy and one pathological. We believe that this approach minimizes the likelihood that the classification model learns patient identities, as the patient’s state remains independent of their identity.
In addition, we exclude recordings labeled 1573-a-n.wav and 87-a-n.wav. The former contains two distinct sound recordings, while the latter is corrupted by an artifact, likely caused by hardware or software errors. Moreover, we exclude several recordings based on the information provided in the comments from the database, stating they contain some artifacts or other problems. Finally, for several recordings, the ”Pathologies” column contains information that the speaker performs the tasks during the recording using their singing voice or that they are a singer, which may lead to incorrect labeling of the subject as pathological. We exclude these samples from the study as well.
Next, we trim remaining recordings to eliminate any potential silent parts. Using the librosa library [41], we trim the leading and trailing parts of the recordings that are 15 dB quieter than the maximum root mean square value of the amplitude in the analyzed recording.
The preprocessing step results in 1636 recordings. The age distribution of the data divided by sex and pathology is described in Table 1 as well as in Figure 4. The complete list of recordings excluded from the experiment, along with the reason for their exclusion, is included in the Supplementary Material. The outcome variable takes the value of 0 if no pathology is diagnosed, and the value of 1 if one or more pathologies are diagnosed.
Table 1: Age distribution among the sex and voice condition in the used data
Female Male Unit
Healthy Pathological Healthy Pathological
Mean 25.38 48.55 31.45 52.38 years
Standard deviation 11.21 15.28 11.52 15.12
Minimum 18.00 18.00 18.00 18.00
25% percentile 20.00 36.00 22.00 41.00
50% percentile 21.00 49.00 28.00 55.50
75% percentile 24.00 60.00 38.00 63.00
Maximum 84.00 94.00 69.00 89.00
Total number of subjects 407 541 252 436 -
Figure 4: Age distribution of healthy and pathological male and female subjects
The whole preprocessing workflow is illustrated in Figure 5.
Figure 5: Preprocessing before feature extraction
3.2 Feature Extraction
Acoustic features we use can be categorized into time domain, spectral, and cepstral features. We rely on various Python libraries for the feature extraction, namely parselmouth [42] for features related to f0 and formants, spkit [43] for Shannon entropy, torchaudio [44] for linear-frequency cepstral coefficients (LFCC), librosa for the remaining acoustic features, and SciPy [45] and NumpPy [46] to calculate statistical values from the extracted features or from the signal. Table 2 describes libraries used to extract each feature. It also contains the list of all features used in this article, along with references to other related papers utilizing these features.
In addition to the aforementioned features, we also included information about the age of the speakers, as age has an effect on the overall quality of the voice and because this data is practically always known.
Table 2: Extracted features, their notation, their use in feature sets, and library used for extraction
Feature Symbol Configuration for feature set creation Python library Used in articles
mean f0 across all window of the signal $\overline{f}_{0}$ used in all feature sets parselmouth [42] [6, 8, 14, 15, 16]
harmonic-to-noise ratio $\overline{HNR}$ used in all feature sets parselmouth [42] [8, 15, 16]
jitter $jitta$ used in all feature sets parselmouth [42] [6, 8, 15, 16]
shimmer $shim$ used in all feature sets parselmouth [42] [6, 8, 15, 16]
age $age$ used in all feature sets -- [15]
standard deviation of f0 $\sigma_{\mathbf{f}_{0}}$ used / not used parselmouth [42], numpy [46] [6, 8]
occurrence of NaN values in f0-related features $NaN$ used / not used -- Novel
pitch difference $\Delta f_{0}$ used / not used parselmouth [42], numpy [46] Novel
Shannon entropy $H$ used / not used spkit [43] [19]
mean values of the first 20 LFCC $\overline{\mathbf{LFCC}}$ used / not used torchaudio [44] [47]
mean values of the first three formants $\overline{\mathbf{f}}$ used / not used parselmouth [42] [23]
skewness $skew$ used / not used SciPy [45] [48]
mean spectral centroid $\overline{S}$ used / not used librosa [41],numpy [46] [49]
mean spectral contrast $\overline{\mathbf{SC}}$ used / not used librosa [41], numpy [46] [49]
mean spectral flatness $\overline{SF}$ used / not used librosa [41], numpy [46] [50]
mean spectral roll-off $\overline{RO}$ used / not used librosa [41], numpy [46] [49]
mean zero-crossing rate $\overline{ZCR}$ used / not used librosa [41], numpy [46] [19]
mean values of the selected MFCC $\overline{\mathbf{MFCC}}$ not used / used 13 / used 20 librosa [41],numpy [46] [5, 9, 14, 16, 17, 18, 20]
mean first derivative of the selected MFCC $\overline{\Delta\mathbf{MFCC}}$ librosa [41], numpy [46] [16, 17, 18, 20]
mean second derivative of the selected MFCC $\overline{\Delta^{2}\mathbf{MFCC}}$ librosa [41], numpy [46] [16, 17, 18, 20]
variance of selected MFCC $\mathbf{\sigma^{2}_{MFCC}}$ used / not used librosa [41],numpy [46] [8, 18]
variance of first derivative of the selected MFCC $\mathbf{\sigma^{2}_{\Delta MFCC}}$ librosa [41], numpy [46] [18]
variance of second derivative of the selected MFCC $\mathbf{\sigma^{2}_{\Delta^{2}MFCC}}$ librosa [41], numpy [46] [18]
3.2.1 Pitch Difference
To improve classification, we introduce a novel feature, the pitch difference. We determine the pitch difference as the difference ($\Delta f_{0}$) between the maximum and minimum values relative to the minimum value of the extracted f0 (Equation 1). Our reasoning is that a healthy voice should maintain a stable frequency for the full duration of the recording compared to the pathological voice, which may fluctuate. Note, that this feature is dimensionless and allows for comparisons across different scales and dataset.
$\Delta f_{0}=\cfrac{\max(\mathbf{f}_{0})-\min(\mathbf{f}_{0})}{\min(\mathbf{f}_{0})}$ (1)
3.2.2 NaN Feature
For eight recordings, specifically 492-a_n.wav, 719-a_n.wav, 720-a_n. wav, 915-a_n.wav, 1338-a_n.wav, 1407-a_n.wav, 1716-a_n.wav, and 2235-a_n.wav, the algorithm extracting the f0, implemented in parselmouth library, fails and therefore, the feature values for $\mu_{f_{0}}$, $\Delta f_{0}$, $\sigma_{f_{0}}$, jitta, and shim are NaN. This is probably caused by the dominance of the disharmonic part for serious cases of voice pathology. We replace the NaN values with zero and add a new binary feature that takes the value of 1 for the occurrence of NaN values. As this inability to detect the fundamental frequency is an important information about the recording, we consider introducing this feature a legitimate approach. Note that these eight recordings represent 2 female and 6 male subjects in our dataset, and all of them suffer from some form of voice pathology according to the database records.
3.2.3 Feature Sets
In the extracted features, there are significant differences in some feature values between male and female patients. This is likely caused by the higher pitch of the female voice compared to the male voice, altering characteristics such as f0 (see Figure 6). Since several features are dependent on f0, even if indirectly, and the patient’s biological sex is known during examination, we treat the data as two separate datasets and train two separate classification models, one for each sex. To verify this approach, we also train models for both sexes together.
We test different configurations of features to see their influence on classification performance. The possible configurations of optional features are listed in Table 2. In total, we generate 20480 feature subsets, each for males, females, and both sexes. Note that exhaustive feature selection would lead to evaluation of more than $10^{49}$ feature sets, therefore, we include f0, HNR, jitter, shimmer and age in all generated feature sets. For MFCC and their derivatives, we also limit the potential combinations to using no coefficients and derivatives, or first 13 and 20 coefficients and their derivatives, respectively. The combinations are then further expanded with the variances of the coefficients and their derivations.
Figure 6: Distribution of the estimated mean f0 values among male and female patients
3.3 Data Augmentation
In order to address the issue of class imbalance in the utilized dataset, which leads to models with high recall and low specificity, we employ the k-means SMOTE algorithm [51]. This technique is specifically applied to the training set to enhance the model’s ability to learn from minority class instances and thus improve its predictive performance.
K-means SMOTE is an advanced oversampling method that combines the clustering capabilities of k-means [52] with the synthetic data generation process of SMOTE [53].
The method has a risk of failing if the locations of the clusters are initialized close to outliers; therefore, we try initializing the method repeatedly, up to a maximum of ten times. If the method fails after ten repetitions, the algorithm is interrupted and SMOTE [53] is used instead.
We apply the SMOTE-based algorithm only to the training data, to prevent data leakage. This approach allows us to mitigate the adverse effects of class imbalance, resulting in a more robust and reliable predictive model.
3.4 Machine Learning Algorithms
We perform binary classification using NB, KNN, DT, random forest (RF), AdaBoost, and SVM. Given the relatively high dimensionality of the feature sets with a limited number of samples, deep learning models could be challenging to apply effectively. Therefore, we believe that using traditional ML algorithms is a suitable approach and aligns with the findings from the existing research.
All evaluated ML algorithms were implemented using the scikit-learn v1.5.2 library [54]. The list of hyperparameters tuned using grid search along with their corresponding values for each algorithm is presented in Table 3. The naming of hyperparameters in the tables in the following text corresponds to the naming of function parameters in scikit-learn.
Table 3: ML algorithm hyperparameters for grid search
Algorithm Hyperparameter Tested values
SVM kernel ”rbf”, ”poly”
degree (”poly” only) 2, 3, 4, 5, 6
gamma 0.5, 0.1, 0.05, 0.01, 0.005, 0.001, ”auto”
C 0.1, 0.5, 1, 5, 10, 50, 100, 500, 1000, 3000, 5000, 7000, 10000, 12000
KNN n_neighbors 1, 3, 5, 7, 9, 11, 13, 15, 17, 19, 21, 23
p 1, 2
weights ”uniform”, ”distance”
DT criterion ”gini”, ”entropy”, ”log_loss”
splitter ”best”, ”random”
min_samples_split 2, 3, 4, 5, 6, 7, 8, 9, 10
max_features ”sqrt”, ”log2”
RF criterion ”gini”
min_samples_split 2, 3, 4, 5, 6
n_estimators 50, 75, 100, 125, 150, 175
max_features ”sqrt”
AdaBoost learning_rate 0.1, 1, 10
n_estimators 50, 100, 150, 200, 250, 300, 350, 400
NB var_smoothing 1e-8, 1e-9
3.5 Validation of Results
It is crucial to establish robust performance metrics to accurately assess model capability in distinguishing between healthy and pathological patients. Equally important is to ensure that the reported validation results are robust, minimizing the influence of randomness to guarantee the reliability and consistency of our conclusions, as well as the possibility to independently reproduce our findings.
3.5.1 Used Metrics
The imbalance between healthy and pathological samples in the dataset (407 healthy females, 541 females with pathologies, 252 healthy males and 436 males with pathologies) introduces bias into commonly used metrics such as accuracy, F1 score, precision, and negative predictive value [55]. Especially, accuracy can dangerously show overoptimistic results and provide misleading information [56].
To address this issue and to provide metrics that reflect class imbalance, we evaluate the model performance using sensitivity (Equation 2), specificity (Equation 3), UAR (Equation 5), geometric mean (GM) (Equation 4), and bookmakers informedness (BM) (Equation 6). As all these metrics provide information about the successful classification, we also evaluate MCC (Equation 7), which, although not entirely unbiased, has a smaller bias compared to accuracy and also takes misclassification into account [55].
True positive (TP) predictions are correctly predicted positive (pathological) samples and true negative (TN) predictions are correctly predicted negative (healthy) samples. False positive (FP) predictions mark positive predictions of negative samples and false negative (FN) predictions mark negative predictions of positive samples.
$\text{Sensitivity}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (2)
$\text{Specificity}=\frac{\text{TN}}{\text{TN}+\text{FP}}$ (3)
$\text{GM}=\sqrt{\text{Sensitivity}\cdot\text{Specificity}}$ (4)
$\text{UAR}=\frac{\text{Sensitivity}+\text{Specificity}}{2}$ (5)
$\text{BM}=\text{Sensitivity}+\text{Specificity}-1$ (6)
$\text{MCC}=\cfrac{\text{TP}\cdot\text{TN}-\text{FP}\cdot\text{FN}}{\sqrt{\begin{aligned} (\text{TP}+\text{FP})\cdot(\text{TP}+\text{FN})\cdot\\ \cdot(\text{TN}+\text{FP})\cdot(\text{TN}+\text{FN})\end{aligned}}}$ (7)
Note that BM, UAR, and GM give the same weight to sensitivity and specificity, regardless of the distribution of positive and negative samples in the dataset. BM and GM penalize the performance on the minority class more compared to UAR, resulting in 0 values for 0 sensitivity (or specificity).
3.5.2 Validation Approach
Due to the limited size of the dataset and its imbalance, we employ stratified 10-fold cross-validation during the grid search, which should lead to less biased results of model performances [57]. Only the training folds are augmented using the k-means SMOTE-based algorithm (see Section 3.3) in each iteration. Then, each feature in the training set is scaled to an interval $[0,1]$ with the min-max scaler. The parameters obtained for the scaling of the training folds are used to scale the validation fold to avoid potential data leakage.
Algorithm 1 Results validation
1:for each ML algorithm do
2:    load results of all models
3:    sort results according to MCC score
4:    take 1000 results with highest MCC score
5:    for each model with corresponding feature set from previous step do
6:         i = 0
7:         while i $<$ 100 do
8:             split the feature set randomly to 10 stratified folds
9:             for each fold do
10:                 use this fold as a validation set and oversample rest of folds with k-means SMOTE algorithm
11:                 find a scaling parameters to scale each feature in training set to interval $[0,1]$
12:                 scale features in validation set using the scaling parameters from previous step
13:                 fit the model
14:                 compute performance metrics
15:             end for
16:             i = i + 1
17:         end while
18:         compute average performance metrics from all repetitions of stratified cross-validation
19:    end for
20:    select the model with corresponding feature set that has the highest average MCC as the best performing model
21:end for
This process yields ten values for each performance metric. We then calculate the mean value of each metric across the folds, following standard cross-validation practices.
The workflow of the single grid search iteration step, from the augmentation to the calculation of the results from the cross-validation, is illustrated in Figure 7.
Figure 7: Single grid search iteration step
After completing the grid search, we identify the top 1000 performing models for each ML algorithm based on MCC. To account for the variance in MCC introduced by the cross-validation splits, we perform 100 times repeated stratified 10-fold cross-validation for these models. This allows us to estimate the average metrics for each classification model and their corresponding standard deviations more reliably. The validation of the best results is shown in Algorithm 1.
Moreover, we determine the top performing feature sets for each ML algorithm. First, we aggregate the results from the grid search to get the average performance as well as the standard deviation for each feature set and ML algorithm combination. Then, we select the best feature set for each ML algorithm based on the highest value of MCC.
3.5.3 Ensuring Reproducibility
To maintain the reproducibility of our findings, we implement multiple mechanisms so that all computations provide the same results and everyone can verify that our results were produced by the code we supply. The main problem arises from the unspecified license of the SVD. Therefore, we cannot share the feature sets, and we implement methods based on SHA256 checksums to validate input data and intermediate results. To make the use of methodology easier, we include a file named svd_information.csv in the code repository, which contains metadata for each recording session, such as age, pathology, etc. The full list of files downloaded for this work is included in the repository along with SHA256 checksums.
Some algorithms used in the experiment were based on randomness, usually derived from the initial random values of trainable parameters. To mitigate these problems, we explicitly initialized pseudo-random number generators (PRNGs) with a seed (value 42 is used). We report exact versions of all software and computing hardware, and strict adherence to our setup is highly recommended.
Most computations are done in floating-point arithmetic, which introduces rounding errors. Moreover, to increase computational speed, software such as compilers does not always fully adhere to exact specifications and reference implementations. To ensure a way for other researchers to reproduce our results, we round some intermediate results where these errors happen, as well as increase floating point precision to minimize these errors.
4 Results
All calculations are implemented in Python 3.12 [58]. The code is available at https://github.com/aailab-uct/Automated-Robust-and-Reproducible-Voice-Pathology-Detection. The ML pipeline and k-means SMOTE-based algorithm are implemented using the imbalance-learn v0.12.4 library [59]. The computations are done on multiple servers all with 2x AMD EPYC 9374F, 64 GB RAM with GNU/Linux OS Ubuntu 24.04 LTS. The libraries used for feature extraction, data augmentation, and model training are listed in the requirements.txt file included in the aforementioned repository.
As we mentioned and explained in Section 3.5, we do not present the accuracy score as the data used for training and evaluation of the dataset is moderately imbalanced. Instead, we provide an alternative in the form of MCC, UAR, GM, and BM. However, if needed, accuracy can still be found in raw results reported in the repository.
4.1 Top Performing Models
The results for the best performing models (see step 20 of Algorithm 1) for females, males and both sexes are presented in Table 4. The corresponding features and hyperparameter setting are in Table 5 for females, in Table 6 for males, and in Table 7 for both sexes. Note, that a zero standard deviation for NB corresponds to a subtle effect of Lidstone smoothing, which is manifested only in higher decimal places.
Table 4: Top model performance for each ML algorithm
Sex Algorithm MCC SEN SPE GM UAR BM
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
F AdaBoost 0.7157 0.0677 0.8942 0.0412 0.8152 0.0583 0.8530 0.0352 0.8547 0.0344 0.7094 0.0687
DT 0.6458 0.0755 0.8330 0.0537 0.8127 0.0627 0.8216 0.0384 0.8229 0.0379 0.6457 0.0758
NB 0.5402 0.0767 0.6905 0.0633 0.8511 0.0547 0.7652 0.0405 0.7708 0.0391 0.5416 0.0781
KNN 0.7063 0.0683 0.8762 0.0448 0.8273 0.0565 0.8506 0.0348 0.8518 0.0344 0.7036 0.0687
RF 0.7155 0.0681 0.8977 0.0411 0.8104 0.0578 0.8521 0.0354 0.8541 0.0345 0.7081 0.0689
SVM 0.7150 0.0670 0.8809 0.0424 0.8313 0.0558 0.8550 0.0342 0.8561 0.0338 0.7122 0.0675
M AdaBoost 0.6375 0.0842 0.7941 0.0575 0.8608 0.0671 0.8255 0.0433 0.8274 0.0431 0.6549 0.0861
DT 0.5241 0.0957 0.7648 0.0640 0.7706 0.0845 0.7656 0.0499 0.7677 0.0491 0.5354 0.0982
NB 0.5438 0.0840 0.6673 0.0669 0.8928 0.0620 0.7703 0.0457 0.7801 0.0439 0.5601 0.0878
KNN 0.6216 0.0821 0.7637 0.0592 0.8778 0.0619 0.8175 0.0426 0.8207 0.0421 0.6414 0.0843
RF 0.6224 0.0912 0.8327 0.0533 0.7968 0.0809 0.8129 0.0479 0.8147 0.0468 0.6294 0.0937
SVM 0.6847 0.0853 0.8530 0.0500 0.8409 0.0736 0.8457 0.0441 0.8469 0.0436 0.6939 0.0873
B AdaBoost 0.6509 0.0587 0.8392 0.0385 0.8154 0.0494 0.8265 0.0299 0.8273 0.0297 0.6545 0.0594
DT 0.5805 0.0627 0.7990 0.0431 0.7865 0.0517 0.7919 0.0319 0.7928 0.0316 0.5855 0.0633
NB 0.5075 0.0578 0.6256 0.0517 0.8839 0.0398 0.7427 0.0329 0.7547 0.0302 0.5095 0.0603
KNN 0.6583 0.0570 0.8120 0.0411 0.8552 0.0429 0.8328 0.0287 0.8336 0.0286 0.6673 0.0572
RF 0.6735 0.0565 0.8641 0.0352 0.8093 0.0489 0.8356 0.0291 0.8367 0.0286 0.6734 0.0572
SVM 0.6925 0.0556 0.8478 0.0366 0.8506 0.0432 0.8487 0.0279 0.8492 0.0278 0.6984 0.0557
Table 5: Top model – feature set configuration for each ML algorithm — females
Algorithm AdaBoost DT NB KNN RF SVM
’learning_rate’:  0.1 ’n_estimators’:  300 ’criterion’:  ’gini’ ’max_features’:  ’sqrt’ ’min_samples_split’:  10 ’splitter’:  ’random’ ’var_smoothing’:  1e-09 ’n_neighbors’:  11 ’p’:  2 ’weights’:  ’distance’ ’criterion’:  ’gini’ ’max_features’:  ’sqrt’ ’min_samples_split’:  4 ’n_estimators’:  175 ’C’:  500 ’gamma’:  0.5 ’kernel’:  ’rbf’
$\overline{f}_{0}$ Y Y Y Y Y Y
$\overline{HNR}$ Y Y Y Y Y Y
$jitta$ Y Y Y Y Y Y
$shim$ Y Y Y Y Y Y
$NaN$ Y Y N N Y N
$age$ Y Y Y Y Y Y
$\sigma_{f_{0}}$ Y N N Y Y Y
$\Delta f_{0}$ Y Y N Y Y Y
$H$ Y N Y N Y N
$\overline{\mathbf{LFCC}}$ Y N Y N N N
$\overline{\mathbf{f}}$ N N N N Y Y
$skew$ Y N N N Y N
$\overline{S}$ Y N N N N N
$\overline{\textbf{SC}}$ Y N N N N N
$\overline{SF}$ N N N Y Y Y
$\overline{RO}$ Y N N N N N
$\overline{ZCR}$ N N Y N Y N
$\overline{\mathbf{MFCC}}$ 20 0 0 0 0 0
$\overline{\Delta\mathbf{MFCC}}$ 20 0 0 0 0 0
$\overline{\Delta^{2}\mathbf{MFCC}}$ 20 0 0 0 0 0
$\mathbf{\sigma^{2}_{MFCC}}$ 20 N N N N N
$\mathbf{\sigma^{2}_{\Delta MFCC}}$ 20 N N N N N
$\mathbf{\sigma^{2}_{\Delta^{2}MFCC}}$ 20 N N N N N
Table 6: Top model – feature set configuration for each ML algorithm — males
Algorithm AdaBoost DT NB KNN RF SVM
’learning_rate’:  0.1 ’n_estimators’:  400 ’criterion’:  ’entropy’ ’max_features’:  ’sqrt’ ’min_samples_split’:  9 ’splitter’:  ’random’ ’var_smoothing’:  1e-09 ’n_neighbors’:  21 ’p’:  2 ’weights’:  ’uniform’ ’criterion’:  ’gini’ ’max_features’:  ’sqrt’ ’min_samples_split’:  3 ’n_estimators’:  175 ’C’:  100 ’gamma’:  0.05 ’kernel’:  ’rbf’
$\overline{f}_{0}$ Y Y Y Y Y Y
$\overline{HNR}$ Y Y Y Y Y Y
$jitta$ Y Y Y Y Y Y
$shim$ Y Y Y Y Y Y
$NaN$ N N N N N N
$age$ Y Y Y Y Y Y
$\sigma_{f_{0}}$ Y N N N Y N
$\Delta f_{0}$ N N N N Y N
$H$ N N N N Y Y
$\overline{\mathbf{LFCC}}$ Y N Y N N Y
$\overline{\mathbf{f}}$ N N Y N Y N
$skew$ Y N Y N N N
$\overline{S}$ Y N N Y Y N
$\overline{\textbf{SC}}$ Y N N N Y Y
$\overline{SF}$ N N N N N Y
$\overline{RO}$ Y N N N Y N
$\overline{ZCR}$ N N N Y N N
$\overline{\mathbf{MFCC}}$ 20 0 0 0 0 0
$\overline{\Delta\mathbf{MFCC}}$ 20 0 0 0 0 0
$\overline{\Delta^{2}\mathbf{MFCC}}$ 20 0 0 0 0 0
$\mathbf{\sigma^{2}_{MFCC}}$ 20 N N N N N
$\mathbf{\sigma^{2}_{\Delta MFCC}}$ 20 N N N N N
$\mathbf{\sigma^{2}_{\Delta^{2}MFCC}}$ 20 N N N N N
Table 7: Top model – feature set configuration for each ML algorithm — both
Algorithm AdaBoost DT NB KNN RF SVM
’learning_rate’:  0.1 ’n_estimators’:  400 ’criterion’:  ’gini’ ’max_features’:  ’sqrt’ ’min_samples_split’:  10 ’splitter’:  ’random’ ’var_smoothing’:  1e-08 ’n_neighbors’:  23 ’p’:  2 ’weights’:  ’uniform’ ’criterion’:  ’gini’ ’max_features’:  ’sqrt’ ’min_samples_split’:  6 ’n_estimators’:  175 ’C’:  5 ’gamma’:  0.5 ’kernel’:  ’rbf’
$\overline{f}_{0}$ Y Y Y Y Y Y
$\overline{HNR}$ Y Y Y Y Y Y
$jitta$ Y Y Y Y Y Y
$shim$ Y Y Y Y Y Y
$NaN$ N Y N N N N
$age$ Y Y Y Y Y Y
$\sigma_{f_{0}}$ Y N N Y Y Y
$\Delta f_{0}$ N N N Y N Y
$H$ N N Y N N Y
$\overline{\mathbf{LFCC}}$ Y N Y N Y Y
$\overline{\mathbf{f}}$ N N N N Y N
$skew$ Y N Y N N N
$\overline{S}$ Y N N Y Y N
$\overline{\textbf{SC}}$ Y N N N N N
$\overline{SF}$ N Y N Y N N
$\overline{RO}$ N N N N N N
$\overline{ZCR}$ N N N Y N Y
$\overline{\mathbf{MFCC}}$ 20 0 0 0 0 0
$\overline{\Delta\mathbf{MFCC}}$ 20 0 0 0 0 0
$\overline{\Delta^{2}\mathbf{MFCC}}$ 20 0 0 0 0 0
$\mathbf{\sigma^{2}_{MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta^{2}MFCC}}$ N N N N N N
Based on the MCC score, the best model for males is SVM. For the female dataset, the best performing model is AdaBoost, while the second best is RF. It is worth mentioning that the SVM model performed in classification of female dataset just slightly worse than both AdaBoost and RF model. For both sexes, the best performing model is SVM. The models for females perform better in general, which might be partially affected by the higher sample size and better ratio between the pathological and healthy samples. The top four female models outperformed the top male models and models including both sexes in MCC.
Our proposed feature, the pitch difference, appears in the feature sets of the best performing models, being more prominent in the female feature sets. Specifically, it is utilized by the best performing female AdaBoost, RF, SVM, and KNN models and by the best performing SVM model for both sexes. Regarding classification of males, the pitch difference was used only in RF model, which was the third best model. The NaN feature was used by the best female models only, namely the AdaBoost and RF models.
Feature sets with 13 MFCC features were not present among the best performing models and 20 MFCC features were used only by the female and male AdaBoost models. The variances of MFCC were used by the male and female AdaBoost models. Notably, the top female AdaBoost models utilized almost complete feature set, where only formants, spectral flatness, and zero-crossing rate were omitted. Additionally, the entropy and LFCC were used by the top female, male, and both-sex models. Generally, all features, except 13 MFCC, were used at least by one top performing model.
4.2 Top Performing Feature Sets
Table 8: Top performing feature set for each ML algorithm — males
Algorithm SVM KNN NB DT RF AdaBoost
Performance metrics
MCC $\mu$ 0.5765 0.5916 0.5440 0.4676 0.6384 0.6031
$\sigma$ 0.0488 0.0391 0.0000 0.0376 0.0109 0.0314
SEN $\mu$ 0.7402 0.7598 0.6467 0.7626 0.8303 0.7730
$\sigma$ 0.0946 0.0085 0.0000 0.0195 0.0053 0.0554
SPE $\mu$ 0.8469 0.8499 0.9125 0.7126 0.8179 0.8441
$\sigma$ 0.0724 0.0416 0.0000 0.0379 0.0096 0.0506
UAR $\mu$ 0.7935 0.8049 0.7796 0.7376 0.8241 0.8085
$\sigma$ 0.0272 0.0210 0.0000 0.0199 0.0057 0.0160
GM $\mu$ 0.7856 0.8021 0.7670 0.7351 0.8233 0.8045
$\sigma$ 0.0345 0.0207 0.0000 0.0206 0.0057 0.0176
BM $\mu$ 0.5871 0.6097 0.5591 0.4753 0.6482 0.6170
$\sigma$ 0.0544 0.0420 0.0000 0.0397 0.0114 0.0320
Features used in feature sets
$\overline{f}_{0}$ Y Y Y Y Y Y
$\overline{HNR}$ Y Y Y Y Y Y
$jitta$ Y Y Y Y Y Y
$shim$ Y Y Y Y Y Y
$NaN$ N Y N N N Y
$age$ Y Y Y Y Y Y
$\sigma_{f_{0}}$ Y N N N N N
$\Delta f_{0}$ Y N N N Y Y
$H$ Y N Y N N N
$\overline{\mathbf{LFCC}}$ N N N N N Y
$\overline{\mathbf{f}}$ Y N N N Y N
$skew$ N N Y N N Y
$\overline{S}$ N Y N N Y N
$\overline{\textbf{SC}}$ Y N N N Y N
$\overline{SF}$ N Y N N N Y
$\overline{RO}$ N N N N N N
$\overline{ZCR}$ N Y N N Y Y
$\overline{\mathbf{MFCC}}$ 20 N N N N 20
$\overline{\Delta\mathbf{MFCC}}$ 20 N N N N 20
$\overline{\Delta^{2}\mathbf{MFCC}}$ 20 N N N N 20
$\mathbf{\sigma^{2}_{MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta^{2}MFCC}}$ N N N N N N
Table 9: Top performing feature set for each ML algorithm — females
Algorithm SVM KNN NB DT RF AdaBoost
Performance metrics
MCC $\mu$ 0.5722 0.6842 0.5391 0.5772 0.7225 0.7196
$\sigma$ 0.0965 0.0391 0.0000 0.0288 0.0074 0.0030
SEN $\mu$ 0.7366 0.8582 0.6710 0.7951 0.9055 0.9070
$\sigma$ 0.1182 0.0227 0.0000 0.0218 0.0050 0.0031
SPE $\mu$ 0.8359 0.8251 0.8670 0.7831 0.8076 0.8030
$\sigma$ 0.0666 0.0173 0.0000 0.0210 0.0028 0.0024
UAR $\mu$ 0.7862 0.8416 0.7690 0.7891 0.8565 0.8550
$\sigma$ 0.0495 0.0191 0.0000 0.0143 0.0035 0.0013
GM $\mu$ 0.7575 0.8407 0.7611 0.7877 0.8544 0.8529
$\sigma$ 0.0880 0.0191 0.0000 0.0145 0.0035 0.0013
BM $\mu$ 0.5724 0.6833 0.5380 0.5782 0.7131 0.7100
$\sigma$ 0.0990 0.0383 0.0000 0.0286 0.0070 0.0027
Features used in feature sets
$\overline{f}_{0}$ Y Y Y Y Y Y
$\overline{HNR}$ Y Y Y Y Y Y
$jitta$ Y Y Y Y Y Y
$shim$ Y Y Y Y Y Y
$NaN$ N Y N Y Y Y
$age$ Y Y Y Y Y Y
$\sigma_{f_{0}}$ Y Y N Y N Y
$\Delta f_{0}$ Y Y N N Y Y
$H$ Y N Y N Y N
$\overline{\mathbf{LFCC}}$ N N N N N N
$\overline{\mathbf{f}}$ N N Y N N N
$skew$ Y N Y N Y N
$\overline{S}$ N N N N Y N
$\overline{\textbf{SC}}$ Y N N N N Y
$\overline{SF}$ N Y N N N N
$\overline{RO}$ Y N N N N N
$\overline{ZCR}$ N N Y N N Y
$\overline{\mathbf{MFCC}}$ 13 N N N N N
$\overline{\Delta\mathbf{MFCC}}$ 13 N N N N N
$\overline{\Delta^{2}\mathbf{MFCC}}$ 13 N N N N N
$\mathbf{\sigma^{2}_{MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta^{2}MFCC}}$ N N N N N N
Table 10: Top performing feature set for each ML algorithm — both
Algorithm SVM KNN NB DT RF AdaBoost
Performance metrics
MCC $\mu$ 0.5936 0.6370 0.5140 0.5198 0.6745 0.6415
$\sigma$ 0.0692 0.0339 0.0000 0.0242 0.0093 0.0175
SEN $\mu$ 0.7514 0.8160 0.6203 0.7793 0.8582 0.8343
$\sigma$ 0.1121 0.0112 0.0000 0.0151 0.0053 0.0099
SPE $\mu$ 0.8428 0.8270 0.8938 0.7438 0.8172 0.8099
$\sigma$ 0.0700 0.0274 0.0000 0.0187 0.0068 0.0078
UAR $\mu$ 0.7971 0.8215 0.7571 0.7616 0.8377 0.8221
$\sigma$ 0.0383 0.0175 0.0000 0.0122 0.0047 0.0083
GM $\mu$ 0.7897 0.8203 0.7434 0.7602 0.8365 0.8210
$\sigma$ 0.0471 0.0174 0.0000 0.0124 0.0047 0.0080
BM $\mu$ 0.5942 0.6430 0.5141 0.5231 0.6754 0.6442
$\sigma$ 0.0765 0.0349 0.0000 0.0244 0.0093 0.0165
Features used in feature sets
$\overline{f}_{0}$ Y Y Y Y Y Y
$\overline{HNR}$ Y Y Y Y Y Y
$jitta$ Y Y Y Y Y Y
$shim$ Y Y Y Y Y Y
$NaN$ N Y N Y N Y
$age$ Y Y Y Y Y Y
$\sigma_{f_{0}}$ N Y N Y Y Y
$\Delta f_{0}$ Y N N N N Y
$H$ Y N Y N Y N
$\overline{\mathbf{LFCC}}$ Y N Y N Y N
$\overline{\mathbf{f}}$ N N N N Y N
$skew$ Y N N N Y Y
$\overline{S}$ N N N N Y Y
$\overline{\textbf{SC}}$ Y N N N N Y
$\overline{SF}$ N N N N Y Y
$\overline{RO}$ Y N N N Y N
$\overline{ZCR}$ N Y Y Y N Y
$\overline{\mathbf{MFCC}}$ 20 N N N N N
$\overline{\Delta\mathbf{MFCC}}$ 20 N N N N N
$\overline{\Delta^{2}\mathbf{MFCC}}$ 20 N N N N N
$\mathbf{\sigma^{2}_{MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta MFCC}}$ N N N N N N
$\mathbf{\sigma^{2}_{\Delta^{2}MFCC}}$ N N N N N N
The best performing feature sets for each ML algorithm are selected by averaging their MCC across all hyperparameter configurations (see Section 3.4). The best performing feature sets in combination with each ML algorithm type are in Table 9 for females, in Tables 8 for males, and in Table 10 for both sexes.
Note that for each ML algorithm, a different number of models was trained due to the different number of possible hyperparameter combinations. Moreover, the training was conducted on a single 10-fold cross-validation split which has a different influence on the final performance of each ML algorithm. Therefore, the performance between models is not comparable.
The results show that the variances of MFCC were absent completely in the best performing feature sets. LFCC were not used in female feature sets and spectral roll-off was not used in male feature sets. The remaining features, including pitch difference and NaN feature, were present at least once.
5 Discussion
Table 11: Estimated mean bias of each metric for the best model for each sex
Metric Female Bias Male Bias Both Bias
Sensitivity 0.8942 0.0000 0.8530 0.0000 0.8478 0.0000
Specificity 0.8152 0.0000 0.8409 0.0000 0.8506 0.0000
Accuracy 0.8603 0.0056 0.8486 0.0016 0.8489 -0.0003
F1-score 0.8796 0.0194 0.8771 0.0293 0.8702 0.0212
Precision 0.8654 0.0367 0.9027 0.0599 0.8938 0.0436
GM 0.8538 0.0000 0.8469 0.0000 0.8492 0.0000
UAR 0.8547 0.0000 0.8469 0.0000 0.8492 0.0000
MCC (normalized) 3 0.8388 -0.0170 0.8378 -0.0091 0.8463 -0.0029
  • 3
    Normalized MCC is transformation of MCC to the interval [0,1]
To support our preference for MCC instead of accuracy, we estimate the biases for each metric for the best model in Table 11, according to [55].
Many studies reported in Related Works section exclude data based on individual pathologies. While this may improve model performance and even allow multimodal classification of individual pathologies, we strongly believe this approach actually limits the applicability. In clinical practice, these models could not be utilized unless the excluded pathologies were also excluded from the possible diagnoses for the examined patients, which may prove impractical for clinical use.
None of the investigated studies reported handling the potential data leakage stemming from training models on datasets containing multiple recordings from the same patients. As there is a possibility that models can learn patterns of individual patients, given enough input, we assume many results of the works in ML-based voice pathology detection might be overestimated due to this error. Therefore, we applied a method to exclude duplicate data based on patient identity. Our proposed approach does not discriminate against either healthy or pathological data and does not lead to overoptimistic results.
Our result is the combination of the best sex-aware models with results weighted by the number of subject of each sex. To our best knowledge, this is the first paper on voice pathology detection combining SVD and ML methods, that is fully reproducible and conforms to the REFORMS practices [60]. We provide the filled in REFORMS checklist in A.
All features, except the three indicated in this paragraph, are obtained from voice recordings and are widely used in the models for voice pathology detection (see subsection Section 2). We consider age legitimate as a feature, as other acoustic features depend on the age of the speaker — for example, changes in speaking fundamental frequency with aging [61]. The pitch difference (see subsection Section 3.2.1) is our proposed feature indicating the change of fundamental frequency during the recording and is extracted in a similar way as other considered features. The NaN feature reflects the fact that it was not possible to estimate the fundamental frequency for the patient (as described in Subsection 3.2.3). As the fundamental frequency is considered one of the dominant features in voice pathology diagnosis, we regard NaN feature as a legitimate approach.
5.1 Limitations
We are aware of several limitations our work is subject to. First, our models were tested using SVD only. The used database does not fully reflect the general population, especially in the proportion of healthy and pathological voices. However, at this time, there is no other suitable database that would reflect the general population better. Despite the justification as the only viable source of data, we cannot extrapolate its performance outside of this dataset. As the data was recorded in a controlled environment, we can assume our models might not be able to perform as well with datasets that are recorded under different conditions.
Moreover, we limit our research only to individuals who are 18 years old and older. Another noteworthy limitation was the available computational capacity, which led to careful decision of the ML algorithms, hyperparameter space, and selected features we drew on throughout our work.
We also acknowledge that many voice and speech disorders extend beyond phonatory function, affecting articulation, prosody, and temporal dynamics, which are best analyzed through continuous speech rather than isolated vowel production. Neuromotoric diseases often present with abnormalities in articulation rate, syllable duration, coarticulation patterns, and prosodic modulation. These features cannot be fully evaluated through sustained vowel phonation alone, as they require an analysis of connected speech, where natural variations in stress, intonation, and fluency become more apparent [62, 63] in the context of the four-level model and its implications for understanding the pathophysiology underlying apraxia of speech and other motor speech disorders [64].
In addition, deficits in speech motor control, imprecise consonant production, and irregularities in rhythm, which are key characteristics of certain neurological disorders, are best observed in spontaneous or structured speech tasks rather than in isolated vowel production. Including continuous speech in voice pathology detection frameworks enables a more comprehensive assessment of speech intelligibility, temporal variability, and segmental articulation, ultimately enhancing diagnostic accuracy and the clinical relevance of machine learning models. Acknowledging these limitations, future research should consider incorporating both sustained phonation and continuous speech samples into pathology detection systems to improve sensitivity to a wider range of voice and speech disorders.
6 Conclusion
In this study, we conducted an extensive comparison of ML-based algorithms which are frequently used in related studies. We determined the best hyperparameter settings and feature combination by the grid search method. Moreover, we tested the influence of sex-based data split on the performance of the algorithms.
Next, we introduced several methodological concepts in the field of voice pathology detection and classification:
  • 1.
    we provided reliable and reproducible results showing the top performance for models based on the selected ML algorithms and hyperparameter settings, reporting realistic performance values,
  • 2.
    we introduced two novel features, pitch difference and NaN feature, which both were represented in the feature subsets that reached the reported best performance,
  • 3.
    we avoided potential introduction of data leakage by appropriate SVD data handling,
  • 4.
    we determined the performance by employing low-biased metrics such as MCC, BM and UAR,
  • 5.
    we presented the top performing feature sets for each ML algorithm with respect to all tested hyperparameter configurations.
Finally, there are several limitations to our work which are mostly based on a lack of suitable databases. Due to this fact, we omitted the external validation of the reached results.
Supplementary information
list_of_excluded_files.csv: recordings excluded along with reasons for exclusion (as described in Section 3.1).
Acknowledgements
Jakub Steinbach and Tomáš Jirsa were supported by their specific university grant IGA A1_FCHI_2024_002.
Declarations
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
All data used are publicly accessible on the following webpage: https://stimmdb.coli.uni-saarland.de
Code availability
The code used to produce all results, along with all supplemental material and information to reproduce our results, is stored in publicly available GitHub repository https://github.com/aailab-uct/Automated-Robust-and-Reproducible-Voice-Pathology-Detection with the following DOI: 10.5281/zenodo.13771573.
Author contribution
Jan Vrba: Conceptualization, Formal analysis, Investigation, Methodology, Project administration, Software, Resources, Supervision, Visualization, Writing – original draft, Writing – review & editing Jakub Steinbach: Conceptualization, Data curation, Formal analysis, Funding acquisition, Investigation, Methodology, Software, Resources, Writing – original draft, Writing – review & editing Tomáš Jirsa: Formal analysis, Data curation, Funding acquisition, Investigation, Methodology, Software, Resources, Validation, Writing – original draft, Writing – review & editing Laura Verde: Conceptualization, Formal analysis, Methodology, Software, Writing – original draft, Writing – review & editing Roberta De Fazio: Formal analysis, Software, Writing – review & editing Zuzana Urbániová: Formal analysis, Data curation, Validation, Writing – review & editing Martin Chovanec: Validation, Writing – review & editing Yuwen Zeng: Validation, Writing – review & editing Kei Ichiji: Validation, Writing – review & editing Lukáš Hájek: Resources, Validation, Writing – review & editing Zuzana Sedláková: Validation, Writing – review & editing Jan Mareš: Funding acquisition, Resources, Supervision, Writing – review & editing Noriyasu Homma: Methodology, Project administration, Supervision, Writing – review & editing
Glossary
BM
bookmakers informedness
CNN
convolutional neural network
DNN
deep neural networks
DT
decision trees
EGG
electroglottogram
f0
fundamental frequency
GFCC
gammatone frequency cepstral coefficients
GM
geometric mean
HNR
harmonic-to-noise ratio
KNN
k-nearest neighbors
LFCC
linear-frequency cepstral coefficients
LPC
linear predictive coefficients
LSTM
long short-term memory
MCC
Matthews correlation coefficient
MFCC
mel-frequency cepstral coefficients
ML
machine learning
MLP
multi-layered perceptron
NB
naive Bayes
NN
neural network
PRNG
pseudo-random number generator
RF
random forest
SMOTE
synthetic minority over-sampling technique
SVD
Saarbrücken Voice Database
SVM
support vector machine
UAR
unweighted average recall
Appendix A Reforms checklist
Module 1: Study design
1a. State the population or distribution about which the scientific claim is made
The population for our scientific claim consists of German females aged 18 to 94 and German males aged 18 to 89. These individuals recorded their voices pronouncing the /a:/ vowel at a normal pitch at the Institute für Phonetik, Universität des Saarlandes (see Table 1 and Figure 4).
1b. Describe the motivation for choosing this population or distribution (1a).
In our research, we examine the feasibility of ML methods for voice pathology detection in adult patients. The SVD is the only available suitable dataset, having a relatively large number of samples, while containing a wide number of diseases, which are also presented in the general population (see Section 2 for reasoning and Section 3.1 for description of the data).
1c. Describe the motivation for the use of ML methods in the study.
We aim to build models exploiting the large number of feature for automatic voice pathology detection system that maximize the Matthews correlation coefficient.
In our research we do not investigate the relationship between feature causality or correlation with the pathology itself (see Section 1).
Module 2: Computational reproducibility
2a. Describe the dataset used for training and evaluating the model and provide a link or DOI to uniquely identify the dataset.
We utilize the publicly available dataset ”Saarbruecken Voice Database” available at: https://stimmdb.coli.uni-saarland.de/help_en.php4. There is no unique identifier and no version control of the database, and we do not have a license to share the dataset, therefore, we provide list of files along with sha256 checksum to ensure reproducibility (see Section 3.1).
2b. Provide details about the code used to train and evaluate the model and produce the results reported in the paper along with link or DOI to uniquely identify the version of the code used.
The code used to produce all results, along with all supplemental material and information to reproduce our results, is stored in publicly available GitHub repository https://github.com/aailab-uct/Automated-Robust-and-Reproducible-Voice-Pathology-Detection with following DOI: 10.5281/zenodo.13771573.
2c. Describe the computing infrastructure used.
All of the code produces the same results regardless of the architecture and operating system of the computer. The results were obtained on multiple servers with 2x AMD EPYC 9374F, 64 GB RAM with GNU/Linux OS Ubuntu 24.04 LTS (see Section 4).
2d. Provide a README file which contains instructions for generating the results using the provided dataset and code.
See the GitHub repository.
2e. Provide a reproduction script to produce all results reported in the paper.
See the GitHub repository.
Module 3: Data quality
3a. Describe source(s) of data, separately for the training and evaluation datasets (if applicable), along with the time when the dataset(s) are collected, the source and process of ground-truth annotations, and other data documentation.
All data are from Saarbruecken Voice Database as described in 2a. As this database is relatively small, we utilize only stratified 10-fold cross-validation without test or evaluation dataset. We downloaded data from SVD on July 12, 2022. The recordings used in our study were recorded between November 20, 1997 and June 16, 2004. The ground truth annotations were obtained by evaluation of stroboscopical recording by database authors [36]. All information related to SVD can be found at https://stimmdb.coli.uni-saarland.de.
3b. State the distribution or set from which the dataset is sampled (i.e., the sampling frame).
As we only adapt the database, there is no information on the methodology of selection for the recording.
3c. Justify why the dataset is useful for the modeling task at hand.
We believe SVD dataset is relevant as it contains various pathologies that are also present in general population. In our study, after removing inappropriate recordings, we worked with 64 various pathologies (see Section 3.1).
3d. State the outcome variable of the model, along with descriptive statistics (split by class for a categorical outcome variable) and its definition.
The outcome variable in this study is the health status of patients, classified into two categories: ’Healthy’ and ’Pathological’. This binary classification is based on information provided by the database, specifically the information about pathologies.
The outcome variable is derived from the table containing patient information. Classification takes the value of 0 if there are no values in the pathology column in the svd_information.csv file, and the value of 1 if one or more pathologies are listed in the column (see Section 3.1). The distribution of healthy / pathological recordings of female and male subjects is provided in the Table 1.
3e. State the sample size and outcome frequencies.
In our study, we utilized 1636 recordings out of 2041 total. In total, 977 are pathological and 659 are healthy. More detailed distribution is explained in Table 1, Figure 4, and Figure 5 (see Section 3.1).
3f. State the percentage of missing data, split by class for a categorical outcome variable.
There was no missing data. However, 1 file (1573-a-n.wav) contains a recording of the time between two sessions and one is corrupted (87-a-n.wav). Moreover, there were multiple recordings marked as corrupt in comments to recording sessions. These recordings were excluded from our research. Finally, there were multiple recordings marked with pathologies ”Gesangsstimme” and ”Sängerstimme”, which probably contain healthy subjects. However, as it is not clear, we exclude this data unless they are diagnosed with another pathology (see Section 3.1).
3g. Justify why the distribution or set from which the dataset is drawn (3b.) is representative of the one about which the scientific claim is being made (1a.).
The used database does not fully reflect the general population, in the sense of proportion of healthy / pathological voices. However, at this time, there is no other suitable database that would reflect the general population better (see Section 6).
Module 4: Data preprocessing
4a. Describe whether any samples are excluded with a rationale for why they are excluded.
From the dataset, we remove 13 corrupted recordings, 40 underage recordings, 17 singers and 335 recordings of patients with multiple recording sessions (except first healthy and first pathological recordings, see Figure 5). See Section 3.1 for rationale and GitHub repository for list of excluded files (misc/ list_of_excluded_files.csv).
4b. Describe how impossible or corrupt samples are dealt with.
When the extraction of f0 was impossible, the feature values for $\overline{f}_{0}$, $\sigma_{\mathbf{f}_{0}}$, $\Delta f_{0}$, jitta, and shim were set to 0 and the binary NaN feature was set to 1 (see Section 3.2).
4c. Describe all transformations of the dataset from its raw form (3a.) to the form used in the model, for instance, treatment of missing data and normalization—preferably through a flow chart.
See Figure 5. More details are in sections 3.1, 3.2, and 3.3.
Module 5: Modeling
5a. Describe, in detail, all models trained.
We utilize multiple ML algorithms for classification (see Table 3) and the k-means SMOTE algorithm for dataset augmentation (see Section 3.3).
5b. Justify the choice of model types implemented.
All ML algorithms are suitable for multi-dimensional data, that we are dealing with (see Section 3.4).
5c. Describe the method for evaluating the model(s) reported in the paper, including details of train-test splits or cross-validation folds.
Information about stratified 10-fold cross-validation and repeated stratified 10-fold cross-validation for the best models is described in the section Section 3.5.
5d. Describe the method for selecting the model(s) reported in the paper.
We performed repeated 10-fold cross-validation to estimate the average value of MCC and its corresponding standard deviation. We select the best model according to this MCC. See Section 3.5.
5e. For the model(s) reported in the paper, specify details about the hyperparameter tuning.
Hyperparameter tuning was approached via the grid search method. The range of hyperparameters was decided after preliminary experiments. See Table 3 for possible hyperparameter values.
5f. Justify that model comparisons are against appropriate baselines.
Our results are comparable to results in Section 2. Regarding reproducibility, we believe we are the first paper combining SVD and ML methods while adhering to the REFORMS checklist. Our research distinguishes from the referred works by not eliminating data based on pathologies, by addressing potential data leakage through duplicities, by not oversampling on full dataset, and avoiding data lekage by improper data scaling. See more explanation in Section 4.
Module 6: Data leakage
6a. Justify that pre-processing (Module 4) and modeling (Module 5) steps only use information from the training dataset (and not the test dataset).
By applying the oversampling algorithm only to the training folds, we aimed to prevent data leakage and ensure that the model’s performance evaluation on the test fold remains unbiased. This approach allowed us to mitigate the adverse effects of class imbalance, resulting in a more robust and reliable predictive model. See Sections 3.3 and 3.5.
The whole process, from preprocessing to validation, is described by Figures 5, 7, and 1.
6b. Describe methods used to address dependencies or duplicates between the training and test datasets (e.g. different samples from the same patients are kept in the same dataset partition).
For patients with multiple recordings of the same type (either all healthy or all pathological), we retained only the oldest recorded sample. For patients with both healthy and pathological recordings, we selected the oldest sample of each type, resulting in a maximum of two recordings per patient — one healthy and one pathological. We believe this approach minimizes the likelihood of the model learning patient identities, as the patient’s classification remains independent of their identity. See Section 3.1.
6c. Justify that each feature or input used in the model is legitimate for the task at hand and does not lead to leakage.
All features, except two, are obtained from voice recordings and are widely used in the models for voice pathology detection (see Section 2). We consider the ”AGE” feature legitimate, as other acoustic features depend on the AGE. I.e. there are changes in speaking fundamental frequency with aging [61].
The feature, that we introduced as ”NaN” reflects the fact, that it was not possible to estimate the fundamental frequency for the patient. As the fundamental frequency is considered as one of the dominant features in voice pathology diagnosis, we consider introducing this ”NaN” feature a legitimate approach. Note, that in total, there are 2 females and 6 males in our dataset, that have NaN value of fundamental frequency, all of them suffer from the voice disorder (see Section 3.2).
Module 7: Metrics and uncertainty
7a. State all metrics used to assess and compare model performance (e.g., accuracy, AUROC etc.). Justify that the metric used to select the final model is suitable for the task.
The choice of metrics, with respect to the class imbalance in the data, is written in the Section 3.5. The claim regarding the best model is based on the Matthews correlation coefficient metric, that is suitable for imbalanced datasets and reflect both successes and errors in the classification.
7b. State uncertainty estimates (e.g., confidence intervals, standard deviations), and give details of how these are calculated.
For each of metrics specified in Section 3.5, we provide also the respective standard deviations that were obtained during the cross-validation procedure which is specified in this section.
7c. Justify the choice of statistical tests (if used) and a check for the assumptions of the statistical test.
We do not use statistical tests in this study.
Module 8: Generalizability and limitations
8a. Describe evidence of external validity.
As we consider SVD database for the only feasible database for our research, it is hard to describe evidence of external validity. See Section 4.
8b. Describe contexts in which the authors do not expect the study’s findings to hold.
First, our model was tested using SVD only. The used database does not fully reflect the general population, in the sense of proportion of healthy / pathological voices. However, at this time, there is no other suitable database that would reflect the general population better. Despite the justification as an only viable source of data, we cannot extrapolate its performance outside of this dataset. Moreover, we limit our research only to individuals that are 18 years old and older. As the data was recorded in a controlled environment, we can assume our models might not be able to perform as well with datasets that are recorded during different conditions. Another noteworthy limitation was the available computational capacity which led to careful decision of the ML algorithms and hyperparameter space we drew from during our work. See Section 4.